solve- if one vertex of an equilateral triangle of side "a" lies at origin and other lies on the line x=(square root of 3)*y the third vertex is..


 

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans: y-axis
Sol:
First vertex lie at the origin O (0, 0). The second vertex A lie on th line
y = \frac{1}{\sqrt{3}}x
Slope of this line is 30 & portion of length ‘a’ of this line is one side of tringle OAB. So the slope of one other side should be 90 as angle b/w sides of OAB should be 60. So the third vertex should lie on this sides which is y-axis. We have
B (\frac{\sqrt{3}a}{2}, \frac{a}{2})
C (0, a)
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty